A gas occupied 100ml at 50 degrees Celsius and 1 atm pressure .the gas is cooled at constant pressure so that volume is reduced to 50ml. What is the final temperature

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
Ideal gas law PV=nRT@constant pressure V is proportional to T Now V1/V2=T1/T2100/50 = (273+50)/T2So T2= 161.5K